OpenStudy (anonymous):

Solve each quadratic equation using factoring?? Pleasee help :) 2x^2+9x-5=0

OpenStudy (theviper):

\(\Large{\color{red}{2x^2+9x-5=0}}\) \(\Large{\color{green}{\implies 2x^2+10x-x-5=0}}\) \(\Large{\color{green}{\implies 2x(x+5)-1(x+5)=0}}\) \(\Large{\color{green}{\implies (2x-1)(x+5)}}\)

OpenStudy (anonymous):

oh lol sorry

OpenStudy (anonymous):

is that gon be your final answer?? like i cant do anything else to it to get x-?=0 nunn like that?